# Enhanced many-body quantum scars from the non-Hermitian Fock skin effect

### v1

@@ Abstract
Figure 4: (a) Schematics of the non-Hermitian Bose-Hubbard model (7) in a tilted optical lattice. The non-reciprocal hopping $J\pm\tilde{\gamma}$ (black arrow) is realized by coupling (blue and violet double arrows) two nearest ground states with an excited state with the laser-induced loss (wavy arrow) on the excited state. $U$ represents the on-site interaction strength, while $\Delta$ refers to the linear potential induced by tilting the optical lattice. Kinetic constraints emerge under $U{=}\Delta\gg J,\tilde{\gamma}$ [32], resulting in the effective non-Hermitian PXP model in Eq.(1). This correspondence is illustrated in the red block that the boson hopping is mapped to spin flipping [66]. (b) Dynamics of population imbalance $I(t)$ for the non-Hermitian Bose-Hubbard model (7), as an experimental indicator for enhanced quantum scars. The initial state is ``$20202020201"$ with $L{=}11$. Other parameters are $V{=}\Delta{=}10J$, and $\tilde{\gamma}/J{=}0.0$ (blue) and $0.3$ (red). The red curve exhibits pronounced and robust revivals, showing the enhancement of quantum scars.
@@
Discussion.—We have shown that asymmetric Fock-state transitions can significantly enhance quantum many-body scarring. Unlike the conventional NHSE, which appears at the physical boundaries of a system, our Fock skin effect leads to many-body state revivals through a delocalized Fock skin accumulation. The latter are enabled by consistent asymmetric transitions across the Hamming layers associated with a particular initial state. We have demonstrated that the dynamics of non-Hermitian scar enhancement can be realized in existing Bose-Hubbard quantum simulators [73, 32]. Furthermore, the rapidly improving quantum computing hardware [76, 77, 78, 79, 80], combined with advancements in quantum algorithms, such as variational circuit optimization [81, 82, 83, 84], non-unitary postselection [85, 81, 83] and advanced error mitigation methods [86, 87], provides another platform for simulating our proposed non-Hermitian PXP and related many-body models  [88, 82, 89, 90].
@@
The mechanisms for stabilizing scar states in disordered systems have recently attracted much attention [91, 92, 93, 94, 95, 96, 97, 98]. The robustness of QMBSs in our model is protected by the Fock skin effect, which is absent in other types of non-Hermitian deformations of the PXP model [99], see SM [66]. Moreover, compared to previous works on disordered scared systems [92], the non-Hermiticity in our case can “shield” all characteristic QMBS signatures, including wavefunction revivals, from disorder. Such robust revivals could also be useful for realizing stable dissipative time crystals under various perturbations [100, 35]. Finally, it would be interesting to explore the consequences of the Fock skin effect introduced here for entanglement phase transitions [101, 102] in various non-Hermitian many-body systems [103, 104, 105, 106, 107, 83, 108].
@@ Acknowledgements.
Acknowledgements.—Exact diagonalization simulations have been performed using QuSpin [109, 110]. F.Q. and C.H.L acknowledges support from the QEP2.0 Grant from the Singapore National Research Foundation (Grant No. NRF2021-QEP2-02-P09) and the Singapore MOE Tier-II Grant (Proposal ID: T2EP50222-0008). J.-Y.D. and Z.P. acknowledge support by the Leverhulme Trust Research Leadership Award RL-2019-015. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No.101034413. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@
[…]
${}^{1}$epartment of Physics, National University of Singapore, Singapore 117542
${}^{2}$chool of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, United Kingdom
${}^{3}$nstitute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria
This supplementary material contains: (SI) a detailed description of the experimental setup for realizing an effective non-Hermitian PXP model using ultracold atoms in an optical lattice; (SII) further characterization of the Fock skin accumulation in three variations of the model with different types of non-Hermiticity; (SIII) analysis of the energy spectrum as a function of non-Hermiticity; (SIV) additional evidence for the enhancement of scarring by the non-Hermiticity in the fidelity dynamics and entanglement entropy of eigenstates.
@@
We consider ${}^{87}$b atoms in a one-dimensional (1D) optical lattice, FIG. S1(a), described by a Bose-Hubbard type Hamiltonian defined in terms of standard boson annihilation and creation operators, $\hat{b}_{j}$ and $\hat{b}_{j}^{\dagger}$:
@@
where $n_{j}=\bra{\psi(t)}\hat{b}_{j}^{\dagger}\hat{b}_{j}\ket{\psi(t)}$ is the occupation density on site $j$. For this quantity, we can drop the global loss $\sum_{j}{\gamma}\hat{b}_{j}^{\dagger}\hat{b}_{j}$ in Eq.(S2[111]. This is justified as any loss that affects all sites similarly does not impact the imbalance.
@@ A.1 Effective non-Hermitian PXP model
To induce kinetic constraints in the Bose-Hubbard model in Eq. (S1), we tune the system to a resonance $U=\Delta\gg J,\tilde{\gamma}$ and assume the total number of bosons is equal to the number of lattice sites, corresponding to filling factor $\nu=N/L=1$. In the absence of atomic loss ($\tilde{\gamma}=0$), Ref. [32] observed that under this resonance condition, the two-body repulsion $U$ and linear potential $\Delta$ interplay in such a way that any triple occupancy on a single site is heavily suppressed, leaving only the transitions of the type“$11$$\leftrightarrow$20202020”, i.e., between two singly occupied adjacent sites and a doubly occupied site. This allows to map the dynamics onto a simpler PXP spin model. Below we will generalize this to the case where atomic loss is present. In terms of experimental implementation, the effective model derived below is expected to hold in the regime $U=\Delta=10J\approx 2.0$ kHz. The tilt potential $\Delta$ can be realized either with gravity or external magnetic fields; for instance, in the setup of Ref. [32] $\Delta_{g}=1632$ Hz from gravity (tilting degree $4^{\circ}$ and $767$ nm spacing for ${}^{87}$b atoms), and $\Delta_{B}=368$ Hz from external magnetic fields, resulting in the quoted value $\Delta=\Delta_{g}+\Delta_{B}\approx 2$ kHz.
Working at the resonance $U=\Delta\gg J,\tilde{\gamma}$, the dynamics governed by the model in Eq. S1 can be describde by the following effective Hamiltonian [32]
$\hat{H}_{\mathrm{eff}}=$ $-\sum_{j=2n}(\left(J+\tilde{\gamma}\right)\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+\left(J-\tilde{\gamma}\right)\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}})$ (S5)
$-\sum_{j=2n+1}(J\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+J\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}}).$
@@
$\hat{H}_{\mathrm{NHPXP}}^{\prime}=\Omega\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\Omega\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$ (S8)
@@
For the quench dynamics, we are primarily interested in the $\left|\mathbb{Z}_{2}\right\rangle=\left|\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow......\right\rangle$ product state of spins, which corresponds to the product state of doublons, $\left|202020......\right\rangle$, in the Bose-Hubbard model. The latter can be experimentally prepared using an optical superlattice [32].

### v2

@@ Abstract
Figure 4: (a) Schematics of the non-Hermitian Bose-Hubbard model (5) in a tilted optical lattice. The non-reciprocal hopping $J\pm\tilde{\gamma}$ (black arrow) is realized by coupling (blue and violet double arrows) two nearest ground states with an excited state with the laser-induced loss (wavy arrow) on the excited state. The on-site interaction $U$ and tilt potential $\Delta$ lead to kinetic constraints under $U=\Delta\gg J,\tilde{\gamma}$ [32]. This results in the effective non-Hermitian PXP model in Eq.(1), and boson hopping is mapped to spin flipping (red block) [67]. (b) Classical simulations of population imbalance $I(t)$ for the non-Hermitian Bose-Hubbard model (5). The initial state is |$\ket{20202020201}$ with $N=L=11$. Other parameters are $V=\Delta=10J$, and $\tilde{\gamma}/J=0.0$ (blue) and $0.3$ (red). The red curve exhibits pronounced revivals due to enhanced quantum scars. (c) Quantum circuit for simulating the model Eq. (6). The non-unitary operation $e^{-itH}$ (pink block) is embedded in an extended unitary (blue block), with physical qubits (green) coupled to the yellow ancilla qubit. The post-selection of $\ket{\uparrow}$ on this ancilla qubit gives the final normalization. (d) Digital simulations of normalized fidelity $F(t)$ using the circuit in (c). We use the IBM Q Brisbane device, with noise conditions shown in SM [67]. The initial state is $\ket{\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow}$. Both numerical results (solid curves) and simulation data (circles) show excellent agreement. Notably, the blue squares ($u=0.4$) demonstrate the robustness of enhanced scars under device noise.
Digital simulation.—Quantum computers emerge as a promising platform due to their great programmability  [77, 78], allowing for the design of sophisticated models. As such, utilizing the IBM quantum processor [79], we implement the digital simulation of the dynamics of the following Ising model under:
$\hat{H}_{V}=V\sum^{L-2}_{j=0}(\mathds{1}-\hat{P}_{j})(\mathds{1}-\hat{P}_{j+1})+\sum_{j\in\mathrm{even}}\hat{X}^{\prime}_{j}+\sum_{j\in\mathrm{odd}}\hat{X}_{j},$ (6)
where a strong interaction $V=10$ is chosen to suppress neighboring $\uparrow$-spins, allowing us to recover the non-Hermitian PXP model. We use non-unitary postselection to embed non-unitary evolution $e^{{}_{i}t\hat{H}_{V}}$ in an extended unitary with an additional ancilla qubit [80, 81, 82], see FIG. 4 (c) and SM [67] for further details. The circuit comprises $8$ physical qubits prepared in the |$\ket{\mathbb{Z}_{2}}$ state and one ancilla qubit, decomposed through variational optimization [81, 83, 82, 84, 85] . The final post-selection on the ancilla $\ket{\uparrow}$ lead to the normalization |$\ket{\psi(t)}=e^{-itH}\ket{\psi_{0}}/\left\|e^{-itH}\ket{\psi_{0}}\right\|$. The signature of scars can be measured through the overlap between |$\ket{\psi(t)}$ and |$\ket{\mathbb{Z}_{2}}$. The simulation results, shown in FIG. 4 (d), demonstrate excellent agreement between exact, noiseless results (solid curves) and noisy simulation data (circles). Importantly, the results in blue ($u=0.4$) highlight the robustness of enhanced scars, showing pronounced revivals throughout this period despite the inherent noise in quantum hardware. Such simulations, without any deliberate effort in error mitigation, demonstrate the capability of the Fock skin effect in improving robustness.
@@
Discussion.—We have shown that asymmetric Fock-state transitions can significantly enhance quantum many-body scarring. Unlike the conventional NHSE, which leads to robust localizations at physical boundaries, our Fock skin effect leads to many-body state revivals with periodic pumping through delocalized Fock skin accumulation. This is enabled by asymmetric transitions across Hamming layers from a specific initial state. Compared to the periodic driving method, the Fock skin effect enables tunable enhancement. We have argued that the non-Hermitian scar enhancement can be realized in existing Bose-Hubbard quantum simulators [74, 32], and further confirmed its robustness with digital simulations on IBMQ hardware.
@@
The mechanisms for stabilizing scar states in disordered systems have recently attracted much attention [86, 87, 88, 89, 90, 91, 92, 93]. The robustness of QMBSs in our model is protected by the Fock skin effect, absent in other non-Hermitian deformations of the PXP model [94], see SM [67]. Moreover, compared to previous works on disordered scarred systems [87], the non-Hermiticity in our case can “shield” all characteristic QMBS signatures, including wavefunction revivals, from disorder. Finally, exploring the impact of Fock skin effect on entanglement phase transitions [95, 96] in various non-Hermitian many-body systems [97, 98, 99, 100, 101, 82, 102] would be also intriguing.
@@ Acknowledgements.
Acknowledgements.—Some of the exact diagonalization simulations have been performed using QuSpin [103, 104]. F.Q. and C.H.L acknowledges support from the QEP2.0 Grant from the Singapore National Research Foundation (Grant No. NRF2021-QEP2-02-P09) and the Singapore MOE Tier-II Grant (Award number: MOE-T2EP50222-0003). J.-Y.D. and Z.P. acknowledge support by the Leverhulme Trust Research Leadership Award RL-2019-015. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No.101034413. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@
Ruizhe Shen1, Fang Qin1, Jean-Yves Desaules2,3, Zlatko Papić2, and Ching Hua Lee1
1Department of Physics, National University of Singapore, Singapore 117542
2School of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, United Kingdom
3Institute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria
This supplementary material contains: (SI) a detailed description of the experimental setup for realizing an effective non-Hermitian PXP model using ultracold atoms in optical lattices; (SII) details of the quantum simulation of the non-Hermitian PXP model on IBM Q device; (SIII) further characterization of the Fock skin accumulation in three variations of the model with different types of non-Hermiticity; (SIV) analysis of the energy spectrum as a function of non-Hermiticity; (SV) additional evidence for the enhancement of scarring by the non-Hermiticity in the fidelity dynamics and entanglement entropy of eigenstates; SVI. statistic behaviors of the non-Hermitian PXP model.
@@
We consider 87Rb atoms in a one-dimensional (1D) optical lattice, FIG. S1(a), described by a Bose-Hubbard type Hamiltonian defined in terms of standard boson annihilation and creation operators, $\hat{b}_{j}$ and $\hat{b}_{j}^{\dagger}$:
@@
where $n_{j}=\bra{\psi(t)}\hat{b}_{j}^{\dagger}\hat{b}_{j}\ket{\psi(t)}/\bra{\psi(t)}\ket{\psi(t)}$ is the occupation density on site $j$. For this quantity, we can drop the global loss $\sum_{j}{\gamma}\hat{b}_{j}^{\dagger}\hat{b}_{j}$ in Eq.(S2[105]. This is justified as any loss that affects all sites similarly does not impact the imbalance.
@@ A.1 Effective non-Hermitian PXP model
To induce kinetic constraints in the Bose-Hubbard model in Eq. (S1), we tune the system to a resonance $U=\Delta\gg J,\tilde{\gamma}$ and assume the total number of bosons is equal to the number of lattice sites, corresponding to filling factor $\nu=N/L=1$. In the absence of atomic loss ($\tilde{\gamma}=0$), Ref. [32] observed that under this resonance condition, the two-body repulsion $U$ and linear potential $\Delta$ interplay in such a way that any triple occupancy on a single site is heavily suppressed, leaving only the transitions of the type“$11$” $\leftrightarrow$ “$20$”, i.e., between two singly occupied adjacent sites and a doubly occupied site. This allows to map the dynamics onto a simpler PXP spin model. Below we will generalize this to the case where atomic loss is present. In terms of experimental implementation, the effective model derived below is expected to hold in the regime $U=\Delta=10J\approx 2.0$ kHz. The tilt potential $\Delta$ can be realized either with gravity or external magnetic fields; for instance, in the setup of Ref. [32] $\Delta_{g}=1632$ Hz from gravity (tilting degree $4^{\circ}$ and $767$ nm spacing for 87Rb atoms), and $\Delta_{B}=368$ Hz from external magnetic fields, resulting in the quoted value $\Delta=\Delta_{g}+\Delta_{B}\approx 2$ kHz.
Working at the resonance $U=\Delta\gg J,\tilde{\gamma}$ and weak dissipation, the dynamics governed by the model in Eq. S1 can be described by the following effective Hamiltonian [32]
$\hat{H}_{\mathrm{eff}}=$ $-\sum_{j\in\text{even}}(\left(J+\tilde{\gamma}\right)\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+\left(J-\tilde{\gamma}\right)\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}})$ (S5)
$-\sum_{j\in\text{odd}}(J\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+J\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}}).$
@@
$\hat{H}_{\mathrm{NHPXP}}^{\prime}=\Omega\sum_{j\in{\rm even}}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\Omega\sum_{j\in{\rm odd}}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$ (S8)
@@
For the quench dynamics, we are primarily interested in the $\left|\mathbb{Z}_{2}\right\rangle=\left|\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow......\right\rangle$ product state of spins, which corresponds to the product state of doublons, $\left|202020......\right\rangle$, in the Bose-Hubbard model. The latter can be experimentally prepared using an optical superlattice [32]. Similarly, the occupation imbalance in the initial $\left|202020......\right\rangle$ state in the Bose-Hubbard model maps to the staggered magnetization in the effective PXP model:
$\bar{Z}(t)=\sum_{i}(-1)^{i}\left\langle\hat{Z}_{i}(t)\right\rangle.$ (S12)
FIG.S2 shows a comparison of the imbalance dynamics in the Bose-Hubbard model (with OBCs) and its effective PXP counterpart (both PBCs and OBCs). Under the weak non-Hermitian condition $\tilde{\gamma}/J=0.3$, both models manifest pronounced revivals (red curves), a signature of enhanced scars. Importantly, these enhanced revivals occur at similar timescales in the natural unit $\hbar/J$.
Figure S2: (a) The imbalance dynamics [Eq. (S4)] for the Bose-Hubbard model described by Eq. (S1). (b) and (c) show the staggered magnetization dynamics [Eq. (S12)] for the effective PXP model described by Eq. (S8), with PBCs (b) and OBCs (c). The initial states are |$\ket{20202020201}$ for (a) and $\ket{\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow}$ for (b) and (c). For all results, we set $J$ ($\Omega=\sqrt{2}J$) as the unit for dynamics. For the two models, the enhanced revivals, induced by the Fock skin effect, emerge under the same condition of $\tilde{\gamma}/J=0.3$, and exhibit similar profiles.
A.2 Coherence time
In our experimental proposal, we set the coupling strength to $J=0.2$ kHz. This results in an observation duration of $t=10/t=50$ ms for the results shown in FIG. S2. Additionally, the non-Hermitian term is given by $\frac{\Omega_{0}^{2}}{\gamma}\approx 0.3J=60$ Hz for Eq. (S3), with $\Omega_{0}=30$ Hz and atom loss $\gamma=15$ Hz.
The cold atom experiment in Ref. [59] suggests the coherence time of $t^{\prime}=1$ ms under the atom loss of $\gamma^{\prime}=1.3$ kHz, which is significantly stronger than our proposed value. Consequently, the condition $\gamma t<\gamma^{\prime}t^{\prime}$ in our setup implies that robust measurements and intrinsic physics of the scar enhancement can be observed in our framework. Generally, 87Rb atoms are known to have a long coherence time, exceeding $100$ ms [106]. Therefore, our proposed time scale of $50$ ms is well within the reasonable range, ensuring that the enhanced scar we predict should be observable.
Appendix B SII. Digital quantum simulation of enhanced scars on the IBM Q processor
Quantum computers can efficiently simulate the dynamics of many-body systems by representing quantum states using qubits [77, 78]. By applying a series of quantum gates, one can then simulate the evolution of complex quantum states. We use the IBM Q platform [79] to demonstrate enhanced scarring on a quantum device. We consider the following model
$\hat{H}_{Q}=V\sum^{L-2}_{j=0}(\mathds{1}-\hat{P}_{j})(\mathds{1}-\hat{P}_{j+1})+\sum_{j\in\mathrm{even}}\hat{X}^{\prime}_{j}+\sum_{j\in\mathrm{odd}}\hat{X}_{j},$ (S13)
where we assume open boundary conditions and focus on the regime $V\gg 1$, which restricts the dynamics to the sector of the Hilbert space described by Eq. (S8). This can be shown formally by applying the Schrieffer-Wolff transformation [107] at the leading order in $1/V$, which results in the effective model $\mathcal{P}\left(\sum_{j\in\mathrm{even}}\hat{X}^{\prime}_{j}+\sum_{j\in\mathrm{odd}}\hat{X}_{j}\right)\mathcal{P}$, where we have defined the global projector $\mathcal{P}=\prod_{j}(\mathds{1}-\ket{\uparrow\uparrow}\bra{\uparrow\uparrow}_{j,j+1})$. The action of $\mathcal{P}$ can be equivalently expressed in terms of local projectors $\hat{P}_{j}$, resulting in the desired non-Hermitian PXP model from the main text. Thus, we expect enhanced scars to be present in the above model (S13), which is easier to implement in terms of native gates on the IBM Q device.
In general, simulating quantum many-body models requires the design of a quantum circuit that represents the Hamiltonian. This setup involves choosing and arranging quantum gates which mimic coupling. However, most native gates on the IBM Q processor are unitary, so we can not directly decompose arbitrary non-unitary operations using these unitary gates. Therefore, to implement the dynamics of the model (S13), we employ the ancilla-based method, demonstrated in Refs. [81, 82]. The idea behind this method is to use additional ancilla qubits that will be measured at the end of the time evolution. While the evolution of the full system (physical and ancilla qubits) is unitary, the evolution of the physical subsystem does not have to be, since it is not a closed system but interacts with the ancilla qubits. The desired time evolution $U(t)=e^{-it\hat{H}_{Q}}$ is embedded in an extended unitary:
$U^{\prime}=\left[\begin{array}[]{cc}zU(t)&B\\ C&D\end{array}\right]\quad\text{as}\quad U^{\prime}=\left[\begin{array}[]{cc}\text{physical}\to\text{physical}&\text{ancilla}\to\text{physical}\\ \text{physical}\to\text{ancilla}&\text{ancilla}\to\text{ancilla}\end{array}\right],$ (S14)
where $z=1/\sqrt{\lambda_{\mathrm{max}}}$, with $\lambda_{\mathrm{max}}$ representing the maximum eigenvalue of $U^{\dagger}(t)U(t)$. This extended unitary $U^{\prime}$ represents the blue block shown in FIG. 4 (c) of the main text, which couples the physical and ancilla qubits. Here, the component $C$ is given by $C=A\sqrt{I-z^{2}\Sigma^{2}}E$, where $A$, $\Sigma$ and $E$ are obtained through the singular value decomposition $U^{\dagger}U=A\Sigma E^{\dagger}$  [80]. The components $B$ and $C$ can then be determined by numerically solving for the QR decomposition
$\left[\begin{array}[]{cc}zU&I\\ C&I\end{array}\right]=U^{\prime}R,$ (S15)
where $R$ is the upper triangular matrix of the QR decomposition. In this process, the block $zU$ and $C$ will not be modified, and the solved $B$ and $D$ can ensure the unitary nature of $U^{\prime}$
We set the first site to be the ancilla qubit, which requires post-selection on $\ket{\uparrow}$, and the rest are for the physical system. In this subspace of the ancilla qubit, this initial state can be expressed as (|$(\ket{\psi_{\rm Physical}},0)^{\mathsf{T}}$, with the ancilla |$\ket{\uparrow}=(1,0)^{\mathsf{T}}$. The extended unitary leads to the result:
[…]
The post-selection of $\ket{\uparrow}$ on the ancilla qubit leads to the normalization as follows
[…]
For measurements on quantum computers, we discard the string with the $\downarrow$ ancilla outcome to achieve the above post-selection. In order to implement this unitary on quantum circuits, we prepare a trainable circuit $M$, as shown in FIG. S3 (a), and minimize the following cost function to precisely replicate the target dynamics
${\bf Cost}=1-|\bra{\psi_{0}}M^{\dagger}U_{\bf tar}\ket{\psi_{0}}|,$ (S18)
[…]
The circuit $M$ is parameterized through $U_{3}$ gates
$U_{3}(\theta,\phi,\lambda)=\left[\begin{array}[]{cc}\cos\left(\frac{\theta}{2}\right)&-e^{i\lambda}\sin\left(\frac{\theta}{2}\right)\\ e^{i\phi}\sin\left(\frac{\theta}{2}\right)&e^{i(\phi+\lambda)}\cos\left(\frac{\theta}{2}\right)\end{array}\right].$ (S19)
The training process involves optimizing the angles $\theta$, $\phi$, and $\lambda$. The trainable circuits for simulations are optimized through the L-BFGS-B method, that can efficiently handle large-scale problems with bounds on the parameter range. The L-BFGS-B method iteratively adjusts the parameters to minimize the cost function towards zero, ensuring that the circuit can reach proper convergence.
Figure S3: (a) The structure of the trainable circuit for Eq. (S18). Each blue block represents a circuit layer. The trainable part is the $U_{3}$ gate described by Eq. (S19). The two-qubit gate is the Echoed Cross-Resonance (ECR) gate. (b) Noise conditions of the IBM Q Brisbane device on $2024/05/08$. We present the error rates of measurement and ECR gates. The simulation in this work was conducted on the chain of 00-$8$ qubits.
In general, deeper trainable circuits can lead to faster convergence in optimization tasks, but they can also be more significantly impacted by noise. Therefore, in designing our trainable circuits, we aimed to reach a balance between achieving efficient convergence and maintaining a manageable circuit depth (i.e., the number of circuit layers).
Through empirical testing, we found that the optimal trainable circuits could be constructed within $8$ layers, shown in FIG. S3 (a). This ensures that the cost function can reach the convergence below $1\%$, and simulations are robust to noise. Each circuit layer is built by 16 $U_{3}$ gates and $8$ ECR gates. The ECR gate is defined by the matrix:
$ECR=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cccc}0&1&0&i\\ 1&0&-i&0\\ 0&i&0&1\\ -i&0&1&0\end{array}\right)$ (S20)
This specific gate is chosen to optimize the noise resilience of circuits. Our simulation results are shown in FIG. 4 (d) (main text), conducted on the IBM Q Brisbane device. For these simulations, we used the chain of 00-$8$ qubits on this device. These results demonstrate that our approach successfully balances the circuit depth with the practical constraints imposed by noise and decoherence, ensuring robust performance on current quantum hardware.